OpenStudy (anonymous):

find the general equation of each line; write equation in the form Ax+By=C slope=2; containing the point (-4, 1)

OpenStudy (anonymous):

Use y = m(x - a) + b (m= slope, (a,b) is given point)

OpenStudy (anonymous):

Slope is always the number in front of x in y= mx + b formula. y = 2x + b Plug in your x, y points to find b, the y-intercept. 1 = 2(-4) + b b = 9 y = 2x +9

OpenStudy (anonymous):

how would i put that in Ax+By=C form?

OpenStudy (anonymous):

move 2x over to the other side: y - 2x = 9

OpenStudy (anonymous):

-2x + y = 9 if you want x and y in that order